a container measures 18cm by 22cm by 30cm. it is 1/2 filled with water. water from a jug is then poured in to the container, increasing the volume of water in the container by 1/3 it's original volume. what was the volume of the water in the container at first??
what was the volume of the water in the container after some water was poured into it?
how much water is needed to fill the container completely??

Answer by Theo(13342)

the container measures 18 cm by 22 cm by 30 cm.
the volume of the container is equal to 18 * 22 * 30 = 11880 cubic cm.
the container is half filled with water.
the volume of the water is therefore equal to 11880 / 2 = 5940 cubic cm.
water from the jug is poured in t its original volume.
the volume of the water is now equal to 5940 + 1/3 * 5940 = 7920 cubic cm.
since the container can hold 11880 cubic cm, then that minus 7920 = 3960 cubic cm of water needed to fill the container completely.
that's what i think.